### v1

@@ 6.3 Multimodal
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, [112] proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
6.4 State-of-the-art Methods for CSLR
This subsection reviews state-of-the-art (SOTA) methods for Continuous Sign Language Recognition (CSLR), focusing on their performance on the three most widely used benchmarks: PHOENIX14 [44], PHOENIX14T [26], and CSL-Daily [209].
As shown in Table 7, the dominant architecture remains the RGB-based Spatial Perception Model+1DCNN+BiLSTM pipeline, where the spatial perception model can be either a 2DCNN or a transformer-based approach. Typically, a ResNet18 backbone extracts spatial features, a 1D CNN or TCN models short-term temporal relations, and a BiLSTM captures long-range context for alignment. VAC-CSLR [122] established a widely adopted training setting to address optimization instability. Many works have since extended this baseline with auxiliary classifiers and additional loss terms to enhance feature learning [83, 4, 14]. Furthermore, the TwoStream-SLR [27] seems to compete worthily the aforementioned pipeline, surpassing it when extra training data are utilized [188] and ranked second in the CSL-Daily [209] benchmark without additional samples.
Recent innovations that improve CSLR include:
  • Data Augmentation Across Datasets: Leveraging visually similar signs from different corpora as training data [188].
  • Modeling Gloss Dependencies: Addressing the independence assumption in vanilla Connectionist Temporal Classification (CTC) loss through conditional modeling [197].
  • Multimodal Inputs: Combining RGB and pose heatmaps to improve robustness [27].
  • Graph-Based Feature Interaction: Using GNNs instead of CNNs to model intra- and inter-frame patch relationships [47].
  • Linguistic Priors: Integrating gloss embeddings or language models to guide learning [58, 57].
  • Pre-trained Vision Backbones: Fine-tuning foundation models like CLIP for richer feature extraction [14].
Table 7: Brief description of current state-of-the-art CSLR methods and their performance on the three most used benchmark (Phoenix 2014 [44], Phoenix 2014 T [26] and CSL-Daily [209]). Methods marked with star (*) have utilized additional training data.
Method Modality Architecture Description Training Setting Dataset WER
SLA-Adapter/LoRA [14] RGB CLIP visual encoder with adopter/LoRA, followed by TCN and BiLSTM VAC-CSLR [122] and early supervision, by adding an extra classifier after the visual encoder. CSL-Daily 25.8%
Phoenix 18.8%/19.3%
Phoenix-T 19.5%/19.4%
Uni-Sign [102]* RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders. Three-stage training, multi-task learning with additional training data CSL-Daily 26%
Swin-MSTP [15] RGB SwinT for spatial feature extraction and a mixture of 1D convolution, convolution with different dilation rates and TLP [73]. VAC-CSLR [122] CSL-Daily 27.1%
Phoenix 18.7%
Phoenix-T 19.7%
SlowFastSign [4] RGB Modified SlowFast with bidirectional feature fusion between the two streams. VAC-CSLR [122] and auxiliary classifiers to provide extra supervision. CSL-Daily 24.9%
Phoenix 18.3%
Phoenix-T 18.7%
MultiSignGraph [47] RGB Multiscale spatial-temporal GNN that considers image patches as the nodes. [47]. VAC-CSLR [122] CSL-Daily 26.4%
Phoenix 19.1%
Phoenix-T 19.1%
CSGC [143] RGB 2DCNN+1DCNN+BiLSTM VAC-CSLR [122], auxiliary losses to maintain inter- and cross-sentence consistency and gloss-prototype. CSL-Daily 26.7%
Phoenix 19%
Phoenix-T 19.5%
CorrNet+ [76] RGB Modification of the CorrNet [74], obtained by adding temporal attention after the correlation and identification modules. VAC-CSLR [122] and TLP [73]. CSL-Daily 28.2%
Phoenix 18.2%
Phoenix-T 19.1%
TCNet [111] RGB Combination of the TLP [73] and the CorrNet [74] VAC-CSLR [122] and an additional regularization term CSL-Daily 29.3%
Phoenix 18.9%
Phoenix-T 19.4%
C2ST [197] RGB SwinT+1DCNN+BiLSTM CGD loss, distillation to maintain the consistency between video- and local-level features and linguistic prior CSL-Daily 25.8%
Phoenix 17.7%
Phoenix-T 18.9%
CoSign-2s [83] Pose Dual stream approach based on ST-GCN, considering both joints and motions. Each stream consists of group-specific branches. Extra supervision via auxiliary classifier and complementary regularization to enhance the robustness to noise. CSL-Daily 27.2%
Phoenix 20.1%
Phoenix-T 20.1%
CorrNet [74] RGB 2DCNN equipped with the proposed identification and correlation modules, followed by 1DCNN+BiLSTM. VAC-CSLR [122] CSL-Daily 30.1%
Phoenix 19.4%
Phoenix-T 20.5%
@@ 6.4 State-of-the-art Methods for CSLR
[188]* RGB, Pose TwoStream-SLR [27] Utilization of extra training data via dictionary construction. Phoenix 16.7%
Phoenix-T 18.5%
@@
TwoStream-SLR [27] RGB, Pose Dual stream (containing S3D blocks) with bidirectional lateral connections. Extra supervision achieved by auxiliary CTC losses and self-distillation applied to additional prediction heads. CSL-Daily 25.3%
Phoenix 18.8%
Phoenix-T 19.3%
@@ 7 Datasets
The performance and generalizability of hand gesture and sign language recognition (HGR/SLR) models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation. Over the years, a wide variety of datasets have been developed to support different HGR tasks, ranging from static and isolated gestures to continuous sign language recognition in real-world settings. These datasets vary widely in terms of the number of gestures or glosses, the types of modalities captured (e.g., RGB, depth, pose, IR), the presence of signer diversity, and the complexity of the sequences. In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition, (ii) Semaphoric Gestures, and (iii) Other Gesture Recognition Tasks. This organization aims to provide a clear overview of the resources available for each task type and facilitate their appropriate use in future research.
@@ 7.1 Isolated and Continuous SLR
The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for Isolated Dynamic SLR (IDSLR) and Continuous SLR (CSLR). The former datasets typically focus on individual signs, recorded in controlled lab environments with a fixed vocabulary, making them well-suited for classification-based models. In contrast, continuous datasets present real-world sequences of signs, often extracted from broadcast material or spontaneous signing, requiring temporal segmentation and sequence modeling approaches. Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcripts.
The RWTH-PHOENIX-Weather 2014 [44] and its extended version PHOENIX-2014-T1 [26] are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. More recent efforts such as BOBSL [12] and OpenASL [160] have significantly expanded the signer base and vocabulary, with hours of British and American Sign Language footage collected from TV and web sources, respectively. Lab-recorded corpora like KETI [90], CSL-Daily [209], and How2Sign [41] incorporate multimodal inputs (e.g., pose, depth) and structured glosses, aiding both CSLR and SLT tasks. These datasets collectively enable the development and benchmarking of models under varying constraints of signer variability, language, modality, and task complexity.
On the other hand, isolated sign datasets such as MS-ASL [85], WLASL [99], and BSL-1K [11] provide extensive vocabularies (up to 2,000 signs) and diverse signers, facilitating the study of vocabulary scalability and signer generalization. Many of these datasets, such as AUTSL [164] and BosphorusSign22k [217], also include pose or depth modalities, enriching multimodal learning. While isolated datasets are primarily used for gesture classification, continuous datasets address the more complex problem of temporal segmentation, gloss prediction, and SLT. A few hybrid datasets like GSL SD/SI [2] and Corpus FinSL [154] provide data that can be used in both isolated and continuous settings. Together, these datasets serve a broad spectrum of research needs, from foundational gesture recognition to high-level translation across multiple sign languages.
@@
Table 8: Overview of the characteristics of each dataset related to Sign Language Understanding (SLU) and Hand Gesture Recognition (HGR). The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.

### v2

@@ 6.3 Multimodal
Beyond sign language, general-purpose CHGR systems have explored multimodality across broader visual sensor inputs. For example, (pretrained_on_jester_3dDenseNet_LSTM) proposes a framework that segments RGB, depth, infrared, and optical flow inputs using a shared U-Net and processes them via separate 3D DenseNet-43 backbones. Extracted features are fused through Canonical Correlation Analysis to align shared representations across modalities.
6.4 Performance of CSLR Methods
Table 4 reports the computational cost and Word Error Rate (WER) of CSLR methods on three benchmark datasets. Leveraging large language models, as in (UNISIGNli2025) and (diffusion-CSLR-geng2026no), yields substantial performance gains without dramatically increasing parameters. SlowFastSign (SlowFastSign) achieves strong accuracy with a moderate number of parameters, while TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) has slightly higher WER, requires twice as many parameters, and involves a more complex training setup, though WER improves with additional training data (Fangyun_Wei_ICCV). C2ST (C2ST) is accurate but contains 187.8M parameters, which may limit real-world deployment. For resource-constrained scenarios, the pose-based CoSign-2s (CoSign) requires only 30.1 GFLOPs, though pose estimation cost is not included. DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) is a lightweight RGB-based model but less accurate. Overall, a trade-off between accuracy and computational cost persists.
Table 4: Comparison of CSLR methods in terms of computational cost and accuracy (Word Error Rate, WER) on three benchmarks: CSL-Daily (Back-Translation-CVPR), Phoenix (PHOENIX-2014), and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ). ”Params” denotes the number of trainable parameters.
Method Year GFLOPs Params CSL-Daily Phoenix Phoenix T
(diffusion-CSLR-geng2026no) 2025 - 86.64M 23.7% 16.8% 17.2%
DSTEN-CSLR (DSTEN-CSLR-yu2025dsten) 2025 366.3 30.19M 28.1% 20.0% 20.3%
SLA-Adapter (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 26.2M 25.8% 18.8% 19.5%
SLA-LoRA (CLIP_SLA_alyami2025clip_CSLR_2025) 2025 - 30.8M 25.8% 19.3% 19.4%
Uni-Sign (UNISIGNli2025) 2025 - - 26% - -
Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) 2025 - - 27.1% 18.7% 19.7%
CORE (CORE-CSLR-zhang2025core) 2025 - - 31.1% 21.7% 21.6%
SD-Net (SD-Net-CSLR-gao2025structure) 2025 - - 27.9% - 20.3%
STNet (STNet-wang2025continuous) 2025 - - 27.2% 19.3% 19.8%
SlowFastSign (SlowFastSign) 2024 - 52.5M 24.9% 18.3% 18.7%
MultiSignGraph (SignGraph_CVPR) 2024 - - 26.4% 19.1% 19.1%
CSGC (CSGC_Rao_Sun_Wang_Wang_Zhang_2024) 2024 - - 26.7% 19% 19.5%
CorrNet+ (corrnetplus_hu2024corrnetsignlanguagerecognition) 2024 - - 28.2% 18.2% 19.1%
TCNet (TCNet_Lu_Salah_Poppe_2024) 2024 932.1 19.3M 29.3% 18.9% 19.4%
C2SLR (C2SLR) 2024 1124.6 32.6M - 20.4% 20.4%
GPGN (GPGN) 2024 - - 30% 20.4% 20.5%
C2ST (C2ST) 2023 1635 187.8M 25.8% 17.7% 18.9%
CoSign-2s (CoSign) 2023 30.1 28.2M 27.2% 20.1% 20.1%
CTCA (CTCA) 2023 - - 29.4% 20.1% 20.3%
CorrNet (CorrNet) 2023 1035.4 20.5M 30.1% 19.4% 20.5%
SEN-CSLR (SEN-CSLR) 2023 1144.2 23.1M 30.7% 21% 20.7%
(Fangyun_Wei_ICCV) 2023 - 105.2M - 16.7% 18.5%
CVT-SLR (CVT-SLR) 2023 - - - 20.1% 20.3%
TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) 2022 - 105.2M 25.3% 18.8% 19.3%
STMC (MULTI-CUE) 2022 1742.1 40.71M - 20.7% 21%
TLP (TemporalLiftPooling) 2022 2950.5 48M - 20.8% 21.2%
SMKD (SMKD) 2021 1032.3 20.3M - 21% 22.4%
VAC-CSLR (VAC-CSLR) 2021 1120.4 22.3M - 22.3% -
7 Summary of SoTA Methods
After reviewing numerous recent works, this section focuses on state-of-the-art (SoTA) methods across tasks, summarizing the field and comparing key methods in terms of architecture, deployment, strengths, and limitations. Table 5 highlights each method’s task, application domain, input modality, architecture, and major strengths and weaknesses.
One insight from the table is that SLU dominates other fields, likely due to its complexity, which requires advanced visual cues (including facial expressions), extensive temporal modeling for dynamic gestures, and classifiers capable of distinguishing highly similar gestures. Consequently, few SoTA methods exist for static SLR, as they struggle to capture hand movements. General-purpose IDHGR methods, used as baselines for tasks like SLU, HRI, and HCI, have also been widely studied; the similarity of gestures in HRI and HCI allows the same methods to serve both domains.
Depth, pose, and other non-RGB modalities are used less frequently and typically in combination with RGB streams. This is because they either cannot represent fine-grained gestures (depth, optical flow) or perform poorly in dynamic recognition (pose estimators often fail on blurry images). RGB-based models—even older or simpler ones—already achieve strong results on SHGR datasets (HAGRIDv2_ARXIV), so adding pose can incur significant computational overhead without substantial accuracy gains. In CSLR, rapid hand movements degrade pose estimator performance, making pose usable only alongside RGB (NEURIPS2022_6cd3ac24_HEATMAPS) and increasing complexity. Thus, pose-based approaches are mostly limited to IDHGR, with some exceptions (CoSign), where incorporating multiple modalities can notably improve performance (SAM_SLR; jiang2021sign-sam-slr-v2), even with late fusion.
Regarding architecture, SHGR’s simplicity allows even small 2DCNNs (EDenseNet; G-CNN-SHARMA2021115657) to achieve very high accuracy ($\geq$95%). However, the limited size and diversity of datasets like (NUS-I-2010; NUS-II-2013; ouhands; triesch2001system) can lead to overfitting, reducing real-world performance. In contrast, IDHGR and especially ISLR rely on more complex models. ST-GCN variants (JOINTS_MOTION_VECTORS_Improved_ST-GCN; TD-GCN) and GCN+Transformer pipelines (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) mainly process skeleton data, often combined with RGB 3DCNNs (e.g., I3D, Res3D) to compensate for pose estimation errors. While many studies focus on improving pose-based streams, RGB modalities are less explored. SAM-SLR (SAM_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) propose advanced separate processing for each modality, representing current SoTA for IDHGR. RGBD models (RAAR3DNet; CTFB_Hampiholi; Yunan_Li) typically use dual-stream 3DCNNs, increasing computational cost. CSLR demands even more complex architectures to capture short-term (gloss-wise) and long-term (global) context. The 2DCNN+1DCNN+BiLSTM pipeline remains common, with relatively fewer parameters than SHGR models (e.g., Gesture-CNN (G-CNN-SHARMA2021115657)  67M vs. large CSLR models  48M (TemporalLiftPooling)). Vision transformers increasingly replace 2DCNN backbones (SWIN_MSTP_Neurocomputing_2025_CSLR; C2ST; CLIP_SLA_alyami2025clip_CSLR_2025) for their attention-based focus on critical regions. TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) employs a dual-stream 3DCNN with multimodal fusion and pose heatmaps (instead of skeleton joints) alongside RGB frames, achieving high accuracy.
As mentioned earlier, SHGR is a relatively easy task, so researchers have not proposed any notable learning strategy. Pose-based methods for ISLR often (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao) insert hand topology prior through SSL. NAS is also utilized to determine the optimal structure of 3DCNNs. Models for CSLR face serious optimization difficulties compared to other methods as a consequence of their complexity. Thus, extra supervision is provided via additional losses attached to various layers of the models (VAC-CSLR; CLIP_SLA_alyami2025clip_CSLR_2025). Leveraging additional training data combined with accurate architectures during training (Fangyun_Wei_ICCV) increases the performance, but makes a straightforward comparison with other methods infeasible.
In summary, LHGR-Net (LHGR-Net) and stand-alone pose-based methods (Yangke_patcog; TD-GCN; CoSign) are suitable for resource-constrained use cases (assuming that the cost of the pose estimation is negligible, which is true especially when the keypoints are extracted by the deployed camera), balancing the computational complexity with the recognition accuracy. TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS), C2ST (C2ST) and (diffusion-CSLR-geng2026no) achieve superior performance for CSLR but come with complex training settings and large computational cost. NLA-SLR (NLA_SLR) and SAM-SLRv2 (jiang2021sign-sam-slr-v2) surpass most of the related methods, but they require substantial computational resources. Uni-Sign (UNISIGNli2025) introduced a novelty, by combining various SLU tasks into a single one. Although it achieves significant performance and unifies a variety of tasks providing a shared framework, the computational complexity makes difficult the real-world deployment and the requirement of training data with detailed annotations prevents the adoption to other sign languages.
Table 5: Comparison of current SoTA methods. They are compared in terms of task, application domain (referred to as ”Domain” for the sake of brevity), input modality (or modalities) and general architecture. The strengths of each method or the key innovations introduced as well as some notable limitations of them, are also mentioned.
Method Task Domain Modality Architecture Strengths or Key Innovation Limitations
FGDSNet (FGDSNet) SHGR HRI RGB Dual-stream hand segmentation model + dual-stream 2DCNN + mid-level fusion The dual-stream classifier addresses any failures of the task-specific segmentation model. The segmentation model and the 2DCNN backbones must be trained separately.
LHGR-Net (LHGR-Net) SHGR HCI RGB Atrous convolutional layers followed by depth separable convolutions and attention. Suitable for resource-constrained deployment. Very small capacity.
2DPSTPP-Net (2DPSTPP) SHGR HCI RGB 2DCNN with custom pooling layers and residual connections Spatial pyramid pooling and residual connections obtain robust representations. The model is relatively large for SHGR.
Gesture-CNN (G-CNN-SHARMA2021115657) SHGR SLU RGB 2DCNN consisting of four convolutional layers The simplicity of the architecture enables easy implementation and training. Relatively large (67M parameters).
SAM-SLR (SAM_SLR), SAM-SLR-v2 (jiang2021sign-sam-slr-v2) IDHGR SLU RGBD, Pose, HHA, OF, DF Six/seven streams followed by late fusion Incorporated a large number of modalities, achieving very high accuracy. Insufficient fusion and computationally expensive.
SignBERT (SignBERT_ICCV), SignBERT+ (SignBERT+) (ISLR) IDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + late fusion The model gains hand topology prior through SSL pretraining (masked hand tokens reconstruction, inspired by BERT) Needs extensive SSL pretraining, the processing and incorporation of the RGB modality may be suboptimal
SignBERT+ (SignBERT+) (CSLR) CDHGR SLU RGB, Pose Spatial GCN + Transformer & 3DCNN + intermediate fusion + BiLSTM
BEST (BEST) IDHGR SLU RGB, Pose dVAE + Spatial GCN + Transformer & 3DCNN + late fusion
NLA-SLR (NLA_SLR) IDHGR SLU RGB, Pose Four-stream network based on S3D with two frame rates and multi-level fusion Novel architecture performing multimodality fusion efficiently. Language-aware label smoothing facilitates training. Robust to visually similar signs recognition. Is affected by the performance of word encoders
Uni-Sign (UNISIGNli2025) IDHGR, CDHGR SLU RGB, Pose A pretrained LLM process features extracted by pose, visual and temporal encoders Provides a unified framework for multiple SLU tasks, increasing the recognition accuracy. Requires additional training data
TD-GCN (TD-GCN) IDHGR - Pose ST-GCN variant Facilitates temporal modeling through non-fixed time-dependent topology Depends on the pose estimation performance.
SBI-DHGR (SBI-DHGR) IDHGR - Pose Multi-stream 1DCNN + BiLSTM followed by late fusion Utilizes features of various levels of granularity. Very large compared to other pose-based methods
DSTSA-GCN (DSTSA-CGN_cui2025dstsa) (4s) IDHGR - Pose Four stream GCN + TCN Group channel-wise and temporal-wise GCNs with dynamic graph topologies improve the performance. The static part of the hand topology is initialized randomly, lacking prior.
SlowFast+NAS (3D_CDC_NAS_TIP_2021) IDHGR - RGBD, OF Multi-stream 3DCNN with 3D-CDC and multi-level fusion Enhanced temporal modeling through central difference convolution and optimal hyperparameter selection through NAS. Needs substantials computational resources (251.4M params, 232.1GFLOPs) compared to similar methods
RAAR3DNet (RAAR3DNet) IDHGR - RGBD Dual stream I3D with adoptable cells, spatial and temporal attention NAS improves the architecture. The spatial attention mechanism is supervised by an off-the-shelf pose estimator. It relies on the correctness of the pose estimator
CTFB (CTFB_Hampiholi) IDHGR - RGBD Dual-stream 3DCNN + Multi-level fusion with transformer-based modules Performs efficient multimodality fusion via multiple attention modules placed in different layers. The scaling to more than two modalities is not straightforward
(diffusion-CSLR-geng2026no) CHGR SLU RGB Visual encoder followed by a Diffusion Model (DM). Using a DM for gloss generation makes it very accurate. Increased complexity.
CLIP-SLA (CLIP_SLA_alyami2025clip_CSLR_2025) CHGR SLU RGB CLIP visual encoder (with MLP adopters or LoRA) + TCN + BiLSTM Early supervision by adding an extra classifier accelerates training, while the utilization of a foundation model enhances the feature extraction. The utilization of CLIP implies high computational cost.
Swin-MSTP (SWIN_MSTP_Neurocomputing_2025_CSLR) CHGR SLU RGB SwinT+1DCNN/TLP Deploying a multiscale TCN, the model is able to recognize continuous signs of varying time durations. Computationally expensive.
C2ST (C2ST) CHGR SLU RGB SwinT+1DCNN+BiLSTM CGD loss (that doesn’t make any independence assumption, compared to CTC) and self-distillation increase the performance. The vision transformer as well as the language model utilized during training increase the computational complexity.
CoSign-2s (CoSign) CHGR SLU Pose ST-GCN variant The computational cost is negligible, thus CoSign-2s perfectly balances efficient execution and accurate recognition. The performance depends on the attributes of the pose estimator.
@@ 7 Summary of SoTA Methods
(Fangyun_Wei_ICCV) CHGR SLU RGB, Pose TwoStream-SLR Utilization of extra training data via dictionary construction. Needs additional training data.
@@
TwoStream-SLR (NEURIPS2022_6cd3ac24_HEATMAPS) CHGR SLU RGB, Pose Dual-stream 3DCNN with bidirectional lateral connections Extra supervision by auxiliary losses and self-distillation applied to additional prediction heads facilitate training. The multi-level fusion incorporates efficiently the two modalities. Increased complexity due to the usage of two streams and an off-the-shelf pose estimator. Very large.
@@ 8 Datasets
The performance and generalizability of gesture recognition models are heavily influenced by the quality, diversity, and scale of the datasets used during training and evaluation. Over the years, a wide variety of datasets have been developed to support different HGR tasks, ranging from static and isolated gestures to continuous sign language recognition in real-world settings. These datasets vary widely in terms of the number of gestures or glosses, the types of modalities captured (e.g., RGB, depth, pose, IR), the presence of signer diversity, and the complexity of the sequences. In the following subsections, we categorize and review the most prominent datasets according to the specific recognition task they target: (i) Isolated and Continuous Sign Language Recognition (IDSLR and CSLR, respectively), (ii) Semaphoric Gestures (both static and dynamic), and (iii) Other Gesture Recognition tasks. This organization aims to provide a clear overview of the resources available for each task type and facilitate their appropriate use in future research.
@@ 8.1 Isolated and Continuous SLR
The datasets that have been developed to support video-based Sign Language Recognition (SLR) can be broadly categorized into those designed for IDSLR and CSLR. The former datasets typically focus on individual signs, recorded in controlled lab environments with a fixed vocabulary, making them well-suited for classification-based models. In contrast, continuous datasets present real-world sequences of signs, often extracted from broadcast material or spontaneous signing, requiring temporal segmentation and sequence modeling approaches. Additionally, some datasets support Sign Language Translation (SLT) by providing sentence-level gloss annotations and parallel spoken language transcriptions.
The RWTH-PHOENIX-Weather 2014 (PHOENIX-2014) and its extended version PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. A semi-automated annotation protocol was proposed, leveraging existing methods for speech recognition and accelerating the process. The signers preform gestures in a constrained environment (i.e. wearing dark clothes in front of a static gray background), probably limiting the generalization of models trained to it. On the contrary, more recent efforts such as BOBSL (BOBSL-BBC) and OpenASL (OpenASL) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively. The more unconstrained environment and the increased number of signers improve the HGR models real-world performance. The lab-recorded KETI (KETI_MDPI) comprises samples captured from two camera angles (i.e. front and side), enlarging diversity. The aim of the dataset is to interpret Korean sign language gestures that are performed in emergency situations using skeleton and RGB data, reducing the need of professionals. Another multimodal dataset collected in a laboratory is the How2Sign (How2Sign-CVPR) for ASL recognition and translation. The entire dataset was captured in the termed Green Screen studio with a single RGB camera and depth sensor, while a subset of it in the Panoptic studio using multiple cameras from various viewpoints, enabling accurate 3D pose estimation. Pose data and text transcripts are also considered in TVB-HKSL-News (TVB-HKSL-News). The samples were collected from TV news with subtitles and Optical Character Recognition was applied to extract the text sequence, while an off-the-shelf pose estimator was utilized to obtain the human keypoints. The RGB-based CSL-Daily (Back-Translation-CVPR) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks. A sign dictionary (referred to as SignDict) was also constructed to be used for other tasks, including sign spotting and segmentation. The Auslan Daily Dataset (Auslan-Daily) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos. These samples may contain multiple persons interacting with each other, which makes the dataset suitable for in-the-wild scenarios.
Other datasets for CSLR do not provide text translations. ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects. HKSL (SignBERT-HKSL) was captured in a laboratory environment, providing RGBD data.
Datasets for IDLSR are larger compared to datasets for CSLR, as they do not require complex annotation making the data collection easier. Many of these datasets, such as CSL (CSL500), GSL iso. (ADADOGLOU), BoshporusSign22k (bosphorussign22k), AUTSL (AUTSL), KArSL (KArSL) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention), include pose or depth modalities, enriching multimodal learning. All of them, except for the AUTSL, were collected in a laboratory with fixed background. Datasets that do not include such modalities, can be extended to them by using off-the-shelf DL pose and depth estimators. The ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) is the only one consisting of pose sequences exclusively. MS-ASL (MS-ASL) and WLASL (WLASL) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity. However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (bsl-1K), CSLD (tim-slr_csld_dataseta) and NCSL (2plus1-D-SLR_and_NCSL). For isolated Irish Sign Language (ISL) recognition, only laboratory datasets have been proposed (e.g. INCLUDE (INCLUDE), IISL2020 (IISL2020_MDPI) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention)), making necessary the collection of an in-the-wild dataset. For Vietnamese SLR, authors in (Multi-VSL-dinh2025sign) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity. Other sign languages have also been considered including Argentinian (e.g. LSA64 (Lsa64argentiniansignlanguage)), Persian (e.g. RKS-PERSIANSIGN (persian_sl_dataset)) and Iranian (e.g. ISLR101 (ISLR101-IRANIAN)). The rarity of some sign languages like them mentioned previously poses obstacles to finding relevant annotators, exacerbating the data scarcity challenge, as it will be discussed later.
@@
Table 6: Overview of the characteristics of each dataset related to SLU. The sign vocabulary (Sign Voc.) corresponds to the number of distinct classes included in each dataset.